Given the following cases for water, designate the state as specifically as possible (e.g., 2-phase and specify further if a saturated vapor or saturated liquid, superheated vapor, compressed liquid) of the water and find the other properties (pressure, temperature, specific volume, specific internal energy, quality) not given. Show your final answers in a table like the one listed below (a) P-1.5 bar, u-3000 kJ/kg (b) T 120 C, v 0.60 m/kg (c) P-3 bar, u 2543.6 kJ/kg (d) T-| 90°C, x-20% (e) P- 25 bar, T-100 C (f) T-240°C, x: 0 u kJ/k 3000 Case P [bar 120 0.60 2543.6 20% 190 100 240 c) 25
Answer #1

Solution in Refere thermodynamic steam table. @ P=1.5 bar, u = 3000 kilka at P=1-5 bar, Ug = 2sig. 7 kJ/xg as u 749 : SuperheV = vg = 0.6058 m3/984 at 3 bat T=190°c, x = 20% as se = 20% or H = 0.20 A 2-phase and sateveated misture at T=190°C, Psat =case PC base) 1.5 1.985 Tec | B (m3/Rg) | CARTIE). state 420.18 | 2.1296 3000 NA , Superheated vapore 120 0.60 1864.03 , 67.1